OpenStudy (anonymous):

A 75.0 kg person steps on a scale in an elevator. The scale reads 761 N. What is the magnitude of the acceleration of the elevator? @studo

OpenStudy (anonymous):

The scale reads \[W = m(g+a)\] in an accelerating elevator, so \[a = W/m - g = 761N/75kg - 9.8 m/s^2 = 0.347 m/s^2\]

OpenStudy (anonymous):

computer is not accepting that as my answer

OpenStudy (anonymous):

Does it specify a value for g? if you use g = 9.81 m/s^2 then the answer becomes 0.337 m/s^2. The process I used is how it's done, though.

OpenStudy (anonymous):

correct!!! thanks

OpenStudy (anonymous):

how did u do it to get 0.3337? computer accepted it

OpenStudy (anonymous):

ok don't worry i now understand what u meant

OpenStudy (osanseviero):

Please dont put @openstudy feedback if it isnt related to this

OpenStudy (anonymous):

ok